I need to factorize this equation

16j^2k-8j^6k^5+60j^3

Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52908)

.

Take  4j^2  out of parentheses.


By the way, it is not an equation. It is an expression.



Answer by Theo(13342)

start with 16j^2k-8j^6k^5+60j^3 *****
factor out a 4 to get:
4*(4j^2k-2j^6k^5+15j^3)
factor out a j^2 to get:
4j^2*(4k-2j^4k^5+15j) *****
this might be acceptable, but you can factor out a 2k from part of the expression in parentheses to get:
4j^2*(2k*(2-j^4k^4)+15j) *****
all three forms should give you the same answer when you randomly assign a value to each of the variables.
i chose 3 for j and 5 for k.
i got:
16j^2k-8j^6k^5+60j^3 = 16*3^2*5-8*3^6*5^5+60*3^3 = -18222660
4j^2*(4k-2j^4k^5+15j) = 4*3^2*(4*5-2*3^4*5^5+15*3) = -18222660
4j^2*(2k*(2-j^4k^4)+15j) = 4*3^2*(2*5*(2-3^4*5^4)+15*3) = -18222660
the first factored form might be acceptable, but, if they are real sticklers, then the second factored form should probably satisfy them.
let me know which one worked for you, if any.